OpenStudy (anonymous):

Sam is observing the velocity of a car at different times. After two hours, the velocity of the car is 54 km/h. After four hours, the velocity of the car is 58 km/h. Part A: Write an equation in two variables in the standard form that can be used to describe the velocity of the car at different times. Show your work and define the variables used. (5 points) Part B: How can you graph the equation obtained in Part A for the first six hours? (5 points)

OpenStudy (anonymous):

Do u get it?

OpenStudy (aum):

Does it say anywhere to assume linear relationship between the velocity of the car and the time? Is this question in the chapter that deals with equations of a straight line?

OpenStudy (anonymous):

Well it depends on what the problem is giving you sometimes it can be a straight line.

OpenStudy (anonymous):

@aum

OpenStudy (anonymous):

@nincompoop Hi could you please help me?

OpenStudy (anonymous):

@FreeWilly Hi

OpenStudy (anonymous):

hello???

OpenStudy (anonymous):

@amistre64 Hi Sorry but could you please help me with this?

OpenStudy (anonymous):

@Ashleyisakitty could u please help me?

OpenStudy (amistre64):

the broken nature of the site is making tags and notifs hard to get thru

OpenStudy (amistre64):

how do we write an equation of a line when we are just given 2 points of reference?

OpenStudy (anonymous):

Thank You finally someone who answered me!

OpenStudy (anonymous):

I have no idea im sorry i dont get how to solve velocity

OpenStudy (amistre64):

its not really about solving velocity, but modeling with the data given

OpenStudy (amistre64):

we are given 2 points of reference, the simplest model would be a straight line between the 2 points: a linear model

OpenStudy (anonymous):

ok

OpenStudy (amistre64):

to do that we need one of the points, and the slope between the points in order to develop a line equation

OpenStudy (amistre64):

we could work a system of equations as well but that may be more advanced then your course calls for

OpenStudy (anonymous):

in the problem it said after 2 hours the cars was 54 km/h and after 4 hours it was 58 km/h so to me it increased by 4 every 2 hours

OpenStudy (amistre64):

thats a good observation, that rate of change is called a slope in a linear equation

OpenStudy (anonymous):

yep

OpenStudy (amistre64):

velocity is equal to, (observed rate of change) times (hrs - a), plus b v = m(t-a) + b but then this isnt in standard form standard form of a line is: Ax + By = C, but we can always rework it into this

OpenStudy (amistre64):

so v = 4/2 (t - 2) + 54 is called point slope format

OpenStudy (amistre64):

4/2 = 2/1 sooo v = 2(t-2)+54 v = 2t -4 +54 -2t +v = 50 but standard form start with a postive so adjust this by a factor of -1 2t -v = -50 would seem fair to me

OpenStudy (amistre64):

graphing it has many approaches as well, the intercept method might be suitable

OpenStudy (anonymous):

ok

OpenStudy (amistre64):

find the point t=0 v=? and plot it find the point t=6 v=? and plot it then draw the line between the 2 points

OpenStudy (anonymous):

ok i get it now Thanks for the GREAT help!